Solve for x:
3sin^2 x = cos^2 x ; 0 < = x < 2x

Answer by Alan3354(69443)

3sin^2 x = cos^2 x ; 0 < = x < 2x (I think you mean < 2pi)
3sin^2(x) = 1 - sin^2(x)
4sin^2 = 1
sin^2 = 1/4
sin(x) = ± 1/2
------------
x = 30, 150, 210, 330 or
x = pi/6, 5pi/6, 7pi/6, 11pi/6 radians